# Dynamic Hologram Generation with Automatic Differentiation

### v1

@@ I Introduction
The holograph computation, as illustrated in Figure 1, is a process to compute the phase pattern ${\bm{\phi}}$ on the SLM plane that can traps at coordinates ${\mathbf{r}}$. The efficient and smooth transition of the trap coordinates is crucial for certain applications, such as arranging neutral atoms into a regular pattern [20]. Bottlenecked by the computational cost of the hologram generation algorithm, the recomputation of the holograms at each step may take too long and cause existing atoms to escape. Recently, neutral network [21] and interpolation based methods [22] are proposed to improve efficiency. In this work, we propose a method with better explainability and better numerical stability. Instead of treating the phase-coordinate relationship as a black box, we treat them as an analytic function ${\bm{\phi}}\to{\mathbf{r}}$ where ${\bm{\phi}}$ is a matrix of phase in the SLM plane and ${\mathbf{r}}$ is a vector of trap coordinates in the image plane. They are related through Fourier optics [23]. Given a velocity field of the trap locations $\mathbf{v}=\frac{\partial{\mathbf{r}}}{\partial t}$, we can derive accurate phase dynamics $\frac{\partial{\bm{\phi}}}{\partial t}$ with automatic differentiation [24]. Technically, we use the weighted Gerchberg-Saxton (WGS) algorithm [18, 25] to compute the holograms, which iterates (inverse) Fourier transformations between these planes to derive the phases. This iterative process determines the relation between ${\bm{\phi}}$ and ${\mathbf{r}}$ implicitly. Hence, we employ the implicit function theorem [26, 27] to derive the phase dynamics $\frac{\partial{\bm{\phi}}}{\partial t}$ concerning $\mathbf{v}$. We emphasize that the process to derive the gradient is computationally efficient, sometimes even faster than the forward computation of the WGS algorithm. Using the phase dynamics, we can continuously evolve the phases for multiple steps to achieve the contiguous evolution of trap locations.
@@ II Method
The goal of dynamic hologram generation is to smoothly transition the traps from initial locations to target locations in the image plane. In this work, we intend to achieve this goal with a computer-generated hologram displayed on SLM. When a laser beam is incident on a certain hologram, light gets diffracted. The wavefront or phase profile of the diffracted beam can be modulated by revising the hologram. The framework of our algorithm for calculating the hologram displayed is shown in Fig. 1. In our algorithm, this process is divided into $n$ steps, each step evolves the trap locations from one keyframe to the next keyframe, where the keyframe is the trap locations at the $j$-th step denoted as ${\mathbf{r}}^{(j)}$, the $i$-th element $r_{i}^{(j)}\in\mathbb{R}^{2}$ is a 2D coordinate, and $k$ is the number of traps. As shown in Fig. 1, at the beginning of the $j$-th step, we are given the $(j-1)$th keyframe in the image plane ${\mathbf{r}}^{(j-1)}$ and the target velocity of the $i$-th trap is defined as $\mathbf{v}^{(j)}_{i}=\frac{r_{i}^{(j)}-r_{i}^{(j-1)}}{\Delta t}$ of these trap locations, where $\Delta t$ is the time interval between two steps. We start by computing the phase generating ${\mathbf{r}}^{(j-1)}$ in the SLM plane as ${\bm{\phi}}^{(j-1)}\in\mathbb{R}^{K\times K}$ using the contiguous-WGS (C-WGS) algorithm, where $K$ is the resolution of the SLM plane. Given the target velocities of the trap locations, we then calculate the phase dynamics $\frac{\partial{\bm{\phi}}^{(j-1)}}{\partial t}$ by automatic differentiation. Finally, we use $\frac{\partial{\bm{\phi}}^{(j-1)}}{\partial t}$ to evolve the phases for multiple steps to achieve the contiguous evolution of trap locations.
@@
Figure 1: (a) The optical setup for hologram generation. The incident light is phase-modulated by SLM, which produces the designed hologram at the Fourier plane of the lens. (b) The framework of gradient-flow weighted Gerchberg-Saxton (WGS) algorithm. $\{r_{i}\}$ is the target trap locations in the focal plane and ${\bm{\phi}}$ is the phase in the SLM plane. (i) Use the WGS with contiguous Fourier transformation to generate the phases on the SLM plane. (ii) Given the target velocities of the target trap locations, obtain the phase dynamics $\frac{d\phi}{dt}$ by using the implicit differentiation of the fixed point. (iii) Evolve the phases for multiple steps for contiguous evolution of trap locations.
@@ II.1 The Contiguous-WGS algorithm
3           $({\bm{\phi}},\mathbf{W})$leftarrow cwgs_step(${\bm{\phi}}$, $\mathbf{W}$, $\mathbf{X}$, $\mathbf{Y}$);
@@ II.2 Gradient-based evolution
To temporally evolve the phases ${\bm{\phi}}$ on the SLM plane, it is imperative to establish the relationship between the phase changing rate $\frac{\partial{\bm{\phi}}}{\partial t}=\frac{\partial{\bm{\phi}}}{\partial{\mathbf{r}}}\frac{\partial{\mathbf{r}}}{\partial t}$. Since the trap velocities $\frac{\partial{\mathbf{r}}}{\partial t}=\mathbf{v}$ is given, the main task is to compute the Jacobian $\frac{\partial{\bm{\phi}}}{\partial{\mathbf{r}}}$. This relationship can be accurately determined through automatic differentiation (AD) [24] of the C-WGS algorithm in Algorithm 1. Given sufficient long iterations, the phase in the SLM plane and the target amplitude converge to fixed points $({\bm{\phi}}^{*},\mathbf{W}^{*})$ that satisfy the following equation:
@@
The derivation of this equation can be found in Appendix A. In practice, we can truncate the series in a certain order to obtain an approximate gradient for stability and efficiency. This Jacobian matrix does not need to be computed explicitly, instead, we use the following relation to compute the disired $\frac{\partial{\bm{\phi}}}{\partial t}$ directly:
@@
Since $\eta(\theta)$ is composed of ${\bm{\phi}}$ and $\mathbf{W}$, we use $\left[\frac{\partial\eta^{*}}{\partial\theta}\right]_{\ {\bm{\phi}}^{*}}$ to label the gradient associated with ${\bm{\phi}}^{*}$. The above equation can be computed straight-forwardly using the trick in Algorithm 2 by utilizing the forward mode AD. In the algorithm, we introduce a dummy variable $\delta t$ in line 4-5 to represent the infinitesimal time interval. It perturbs the coordinates and this perturbation is reflected in the DFT matrices $\mathbf{X}$ and $\mathbf{Y}$ at lines 6-7. Then the program runs the cwgs_step function for $m$ times. Since ${\bm{\phi}}$ and $\mathbf{W}$ are already fixed points, they are not changed during the iteration (lines 8-10). This seemingly trivial computation becomes powerful when combined with the forward mode AD (line 13). We set the dummy variable $\delta t$ to 00. It has zero effect to the output, but carries gradient. The AD engine associates each variable with a gradient field, and updates the gradients as the computation goes. Along with the output, the gradient information $\frac{\partial{\bm{\phi}}}{\partial t}$ is also obtained. Forward mode AD is efficient in obtaining the gradient of multiple output variables with respect to single input variables, which is suitable for our problem. AD engines, such as ForwardDiff.jl [32] in Julia and PyTorch [33] in Python, can perform forward mode AD and obtain accurate gradients, while only introduce a constant overhead. One can verify that the gradient corresponds to Equation 3 and  Equation 2 with the series expansion truncated to the $m$-th order. The finite order error is analyzed numerically as shown in Figs. 13 and 14.
@@ 4           $\mathbf{Y}_{vy}\leftarrow e^{-2\pi ivy}$;
6                $({\bm{\phi}},\mathbf{W})$leftarrow cwgs_step(${\bm{\phi}}$, $\mathbf{W}$, $\mathbf{X}$, $\mathbf{Y}$);
@@ III.1 Movement in One Direction
As shown in previous example, the displacement distance $\Delta x$ is directly proportional to $\Delta\phi$, enabling the WGS algorithm to attain the exact solution. However, in situations with more complex patterns, the WGS algorithm can only offer phase approximations within the SLM plane. As a result, discrepancies in phases may arise even with small movements. An example demonstrating this is presented in Appendix B
@@ III.2 Squeezing along one direction
After a single round of WGS calculation, we can derive the gradient $\frac{d\phi}{dt}$ of fixed points according to the second line of Eq. 11 instead of solving the linear equation in the first line for greater stability and efficiency. Then this gradient is applied in the phase evolution process $\phi$ in the SLM plane. However, with increasing time or distance applied to the gradient, the amplitude $B$ of the image in the focal plane will gradually decay, as shown in Fig. 5(a). To quantitatively analyze the decay phenomena, we plot the maximum amplitude $B_{\mathrm{max}}$ of the points influenced with moving distance $\Delta x$ per pixel of the center along the six steps in Fig. 5(a). We find that the amplitude decays rapidly from 0.7 to 0.2 after 0.5 to 1 pixel movement (the initial amplitude is 1), which is depicted in Fig. 5 (b) and (c). The long-distance result is shown in the red-box point in Appendix C
@@ III.4 Rearrangement of traps
Rearranging the position of traps is crucial for atom loading. In this example, we first generate a randomly positioned about half-filled $20\times 20$ grid (256 points by randomly picked) to a full-filled $16\times 16$ grid in square pattern (shown in Fig. 8). We employ the Hungarian algorithm [35] to match the points in the two configurations, ensuring that the total displacement is minimized. Compared with methods of moving from outside to get a full-filled pattern [36], this manipulation is an efficient alternative way to obtain the desired pattern in experiments [37]. We also calculated the displacement $\Delta x$ per pixel of the central points and the maximum amplitude $B_{\mathrm{max}}$ over successive steps (depicted in Fig. 9) to demonstrate the feasibility of the current methodology.
@@ IV Experiment
To benchmark our method’s performance, we experimentally simulated trap rearrangement using a LCOS-SLM (Hamamatsu, X15213-02L). The SLM has a resolution of $1272\times 1024$ and a frame rate of 60 Hz, limited by the DVI transmission rate. We randomly generated initial trap positions on an 18×18 square grid with a filling factor of 0.44. We then calculated the phase using our method and projected it onto the SLM in real time to rearrange the traps into a defect-free $12\times 12$ configuration. Throughout each movement, we monitored the positions and intensities of optical tweezers using a CMOS camera in the Fourier plane. Fig. 10 shows the displacement $\Delta x$ and maximum intensity $B_{\mathrm{max}}$ of one traced optical trap.
We accelerated phase generation on an NVIDIA RTX 4090 GPU. Phase generation takes 75 ms per 10 frames, including 20 iterations of WGS calculation and gradient evolution. Projection requires 128 ms per 10 frames, which aligns with the standard frame rate but is not fast enough compared to the phase generation.
@@ V Discussion
In summary, integrating automatic differentiation with the weighted Gerchberg-Saxton (WGS) algorithm enables seamless manipulation of optical focus arrays. The numerical analysis demonstrates that, in the case of single-point movement, the gradient can accurately adjust the trap locations by any desired distance. In other scenarios, the gradient can accurately adjust the image by at least 0.5 pixels without significant amplitude decay, making it suitable for various pattern movements. While 0.5 pixels may not represent a large displacement, it is important to note that the gradient computation is highly efficient, requiring even less time than the hologram computation itself. By leveraging the gradient, we can enhance the continuity of trap coordinate changes, which may help reduce atom loss during the loading process in realistic atom array experiments. Compared to previous methods, a significant advantage of our approach is its explainability. The gradient can be easily understood and analyzed, which is crucial for experimentalists to comprehend why the method succeeds or fails.
@@ VI Acknowledgment
We acknowledge the very helpful discussion with Wenlan Chen. This project is partially supported by the National Natural Science Foundation of China under grant nos. 92270107, 12404568, T2225018, and T2121001.
@@ Appendix A Implicit function theorem
where we use $[\cdot]_{\bm{\phi}}$ to label the part associated with ${\bm{\phi}}$ and the corresponding linear coefficients in Eq. 9 and  Eq. 10 should be revised as:
@@ Appendix B Four points movement
Figure 13: The phase difference $\Delta\phi_{\mathrm{implicit}}$ obtained from the solution of the linear equation and $\Delta\phi_{\mathrm{expand}}$ from the series expansion in Eq. 11. The label $-n$ denotes the number of series used.

### v2

@@ I Introduction
The hologram computation, as illustrated in Figure 1, is a process to compute the phase pattern $\bm{\phi}$} on the SLM plane that creates traps at coordinates $\mathbf{r}$}. The efficient and smooth transition of the trap coordinates is crucial for certain applications, such as arranging neutral atoms into a regular pattern [20]. Bottlenecked by the computational cost of the hologram generation algorithm, the recomputation of the holograms at each step may take too long and cause existing atoms to escape. Recently, neural network [21] and interpolation based methods [22] are proposed to improve efficiency. In this work, we propose a method with better explainability and better numerical stability. Instead of treating the phase-coordinate relationship as a black box, we treat them as an analytic function ${\bm{\phi}}\to{\mathbf{r}}$ where $\bm{\phi}$} is a matrix of phase in the SLM plane and $\mathbf{r}$} is a vector of trap coordinates in the image plane. They are related through Fourier optics [23]. Given a velocity field of the trap locations $\mathbf{v}=\frac{\partial{\mathbf{r}}}{\partial t}$, we can derive accurate phase dynamics $\frac{\partial{\bm{\phi}}}{\partial t}$ with automatic differentiation [24]. Technically, we use the weighted Gerchberg-Saxton (WGS) algorithm [18, 25] to compute the holograms, which iterates (inverse) Fourier transformations between these planes to derive the phases. This iterative process determines the relation between $\bm{\phi}$} and $\mathbf{r}$} implicitly. Hence, we employ the implicit function theorem [26, 27] to derive the phase dynamics $\frac{\partial{\bm{\phi}}}{\partial t}$ concerning $\mathbf{v}$. We emphasize that the process to derive the gradient is computationally efficient, sometimes even faster than the forward computation of the WGS algorithm. Using the phase dynamics, we can continuously evolve the phases for multiple steps to achieve the contiguous evolution of trap locations.
@@ II Method
The goal of dynamic hologram generation is to smoothly transition the traps from initial locations to target locations in the image plane. In this work, we intend to achieve this goal with a computer-generated hologram displayed on SLM. When a laser beam is incident on a certain hologram, light gets diffracted. The wavefront or phase profile of the diffracted beam can be modulated by revising the hologram. The framework of our algorithm for calculating the hologram displayed is shown in Fig. 1. In our algorithm, this process is divided into $n$ steps, each step evolves the trap locations from one keyframe to the next keyframe, where the keyframe is the trap locations at the $j$-th step denoted as ${\mathbf{r}}^{(j)}$, the $i$-th element $r_{i}^{(j)}\in\mathbb{R}^{2}$ is a 2D coordinate, and $k$ is the number of traps. As shown in Fig. 1, at the beginning of the $j$-th step, we are given the (j1)(j-1) -th keyframe in the image plane ${\mathbf{r}}^{(j-1)}$ and the target velocity of the $i$-th trap is defined as $\mathbf{v}^{(j)}_{i}=\frac{r_{i}^{(j)}-r_{i}^{(j-1)}}{\Delta t}$ of these trap locations, where $\Delta t$ is the time interval between two steps. We start by computing the phase generating ${\mathbf{r}}^{(j-1)}$ in the SLM plane as ${\bm{\phi}}^{(j-1)}\in\mathbb{R}^{K\times K}$ using the contiguous-WGS (C-WGS) algorithm, where $K$ is the resolution of the SLM plane. Given the target velocities of the trap locations, we then calculate the phase dynamics $\frac{\partial{\bm{\phi}}^{(j-1)}}{\partial t}$ by automatic differentiation. Finally, we use $\frac{\partial{\bm{\phi}}^{(j-1)}}{\partial t}$ to evolve the phases for multiple steps to achieve the contiguous evolution of trap locations.
@@
Figure 1: (a) The optical setup for hologram generation. The incident light is phase-modulated by the SLM, which produces the designed hologram at the Fourier plane of the lens. (b) The framework of gradient-flow weighted Gerchberg-Saxton (WGS) algorithm. $\{r_{i}\}$ is the target trap locations in the focal plane and $\bm{\phi}$} is the phase in the SLM plane. (i) Use the WGS with contiguous Fourier transformation to generate the phases on the SLM plane. (ii) Given the target velocities of the target trap locations, obtain the phase dynamics $\frac{d\phi}{dt}$ by using the implicit differentiation of the fixed point. (iii) Evolve the phases for multiple steps for contiguous evolution of trap locations.
@@ II.1 The Contiguous-WGS algorithm
3     $({\bm{\phi}},\mathbf{W})$ $\leftarrow$ cwgs_step($\bm{\phi}$}, $\mathbf{W}$, $\mathbf{X}$, $\mathbf{Y}$);
@@ II.2 Gradient-based evolution
To temporally evolve the phases $\bm{\phi}$} on the SLM plane, it is imperative to establish the relationship between the phase changing rate $\frac{\partial{\bm{\phi}}}{\partial t}=\frac{\partial{\bm{\phi}}}{\partial{\mathbf{r}}}\frac{\partial{\mathbf{r}}}{\partial t}$. Since the trap velocities $\frac{\partial{\mathbf{r}}}{\partial t}=\mathbf{v}$ are given, the main task is to compute the Jacobian $\frac{\partial{\bm{\phi}}}{\partial{\mathbf{r}}}$. This relationship can be accurately determined through automatic differentiation (AD) [24] of the C-WGS algorithm in Algorithm 1. Given sufficient long iterations, the phase in the SLM plane and the target amplitude converge to fixed points $({\bm{\phi}}^{*},\mathbf{W}^{*})$ that satisfy the following equation:
@@
The derivation of this equation can be found in Appendix A. Intuitively, any fixed point can be viewed as the result of iterating a step infinitely many times. Due to the chain rule, each iteration contributes to the gradient, and the total gradient is the sum of these contributions—this is precisely what the series expansion represents in Eq. 2. In practice, we can truncate the series in a certain order to obtain an approximate gradient for stability and efficiency. This Jacobian matrix does not need to be computed explicitly, instead, we use the following relation to compute the disired $\frac{\partial{\bm{\phi}}}{\partial t}$ directly:
@@
Since $\eta(\theta)$ is composed of $\bm{\phi}$} and $\mathbf{W}$, we use $\left[\frac{\partial\eta^{*}}{\partial\theta}\right]_{\ {\bm{\phi}}^{*}}$ to label the gradient associated with $\bm{\phi}$$^{*}$. The above equation can be computed straight-forwardly using the trick in Algorithm 2 by utilizing the forward mode AD. In the algorithm, we introduce a dummy variable $\delta t$ in line 4-5 to represent the infinitesimal time interval. It perturbs the coordinates and this perturbation is reflected in the DFT matrices $\mathbf{X}$ and $\mathbf{Y}$ at lines 6-7. Then the program runs the cwgs_step function for $m$ times. Since $\bm{\phi}$} and $\mathbf{W}$ are already fixed points, they are not changed during the iteration (lines 8-10). This seemingly trivial computation becomes powerful when combined with the forward mode AD (line 13). We set the dummy variable $\delta t$ to 0. It has zero effect to the output, but carries gradient. The AD engine associates each variable with a gradient field, and updates the gradients as the computation goes. Along with the output, the gradient information $\frac{\partial{\bm{\phi}}}{\partial t}$ is also obtained. Forward mode AD is efficient in obtaining the gradient of multiple output variables with respect to single input variables, which is suitable for our problem. AD engines, such as ForwardDiff.jl [32] in Julia and PyTorch [33] in Python, can perform forward mode AD and obtain accurate gradients, while only introduce a constant overhead. One can verify that the gradient corresponds to Equation 3 and  Equation 2 with the series expansion truncated to the $m$-th order. The finite order error is analyzed numerically as shown in Figs. 13 and 14.
@@ 4     $\mathbf{Y}_{vy}\leftarrow e^{-2\pi ivy}$;
6       $({\bm{\phi}},\mathbf{W})$ $\leftarrow$ cwgs_step($\bm{\phi}$}, $\mathbf{W}$, $\mathbf{X}$, $\mathbf{Y}$);
@@ III.1 Movement in One Direction
As shown in the previous example, the displacement distance $\Delta x$ is directly proportional to $\Delta\phi$, enabling the WGS algorithm to attain the exact solution. However, in situations with more complex patterns, the WGS algorithm can only offer phase approximations within the SLM plane. As a result, discrepancies in phases may arise even with small movements. An example demonstrating this is presented in Appendix B
@@ III.2 Squeezing along one direction
After a single round of WGS calculation, we can derive the gradient $\frac{d\phi}{dt}$ of fixed points according to the second line of Eq. 11 instead of solving the linear equation in the first line for greater stability and efficiency. Then this gradient is applied in the phase evolution process $\phi$ in the SLM plane. However, with increasing time or distance applied to the gradient, the amplitude $B$ of the image in the focal plane will gradually decay, as shown in Fig. 5(a). To quantitatively analyze the decay phenomena, we plot the maximum amplitude $B_{\mathrm{max}}$ of the points influenced with moving distance $\Delta x$ per pixel of the center along the six steps in Fig. 5(a). We find that the amplitude decays rapidly from 0.7 to 0.2 after 0.5 to 1 pixel movement (the initial amplitude is 1), which is depicted in Fig. 5 (b) and (c). Therefore, we recommend recalculating the keyframes after a 0.5-pixel movement to prevent significant amplitude decay. The long-distance result is shown in the red-box point in Appendix C
@@ III.4 Rearrangement of traps
Rearranging the position of traps is crucial for atom loading. In this example, we first generate a randomly positioned about half-filled $20\times 20$ grid (256 points by randomly picked) to a full-filled $16\times 16$ grid in square pattern (shown in Fig. 8). We employ the Hungarian algorithm [35] to match the points in the two configurations, ensuring that the total displacement is minimized. Compared with methods of moving from outside to get a fully filled pattern [36], this manipulation is an efficient alternative way to obtain the desired pattern in experiments [37]. We also calculated the displacement $\Delta x$ per pixel of the central points and the maximum amplitude $B_{\mathrm{max}}$ over successive steps (depicted in Fig. 9) to demonstrate the feasibility of the current methodology.
@@ IV Experiment
To benchmark our method’s performance, we experimentally simulated trap rearrangement using a LCOS-SLM (Hamamatsu, X15213-02L). The SLM has a resolution of $1272\times 1024$ and a frame rate of 60 Hz, limited by the DVI transmission rate. We randomly generated initial trap positions on an 18×18 square grid with a filling factor of 0.44. We then calculated the phase using our method and projected it onto the SLM in real time to rearrange the traps into a defect-free $12\times 12$ configuration. We accelerated phase generation on an NVIDIA RTX 4090 GPU. Phase generation takes 75 ms per 10 frames, including 20 iterations of WGS calculation and gradient evolution. Projection requires 128 ms per 10 frames, which aligns with the standard frame rate but is not fast enough compared to the phase generation.
Throughout each movement, we monitored the positions and intensities of optical tweezers using a CMOS camera in the Fourier plane. Fig. 10 shows the displacement $\Delta x$ and maximum intensity $B_{\mathrm{max}}$ of one traced optical trap. Using realistic experimental intensity fluctuations, we simulated the decay of atoms in optical traps similar to [21], more informations see Appendix E. The results demonstrate that the optical traps can be rearranged without significant intensity decay, indicating the effectiveness of our method for real-time optical trap manipulation.
@@ V Discussion
In summary, integrating automatic differentiation with the weighted Gerchberg-Saxton (WGS) algorithm enables seamless manipulation of optical focus arrays. The numerical analysis demonstrates that, in the case of single-point movement, the gradient can accurately adjust the trap locations by any desired distance. In other scenarios, the gradient can accurately adjust the image by at least 0.5 pixels without significant amplitude decay, making it suitable for various pattern movements. While 0.5 pixels may not represent a large displacement, it is important to note that the gradient computation is highly efficient, requiring even less time than the hologram computation itself. By leveraging the gradient, we can enhance the continuity of trap coordinate changes, which may help reduce atom loss during the loading process in realistic atom array experiments. Compared to previous methods, a key advantage of our approach is its explainability, along with its easy extensibility to 3D GS algorithms or any other fixed-point-iteration-based method. The gradient can be easily understood and analyzed, which is crucial for experimentalists to comprehend why the method succeeds or fails.
@@ VI Acknowledgment
We acknowledge the very helpful discussion with Wenlan Chen. This project is partially supported by the National Natural Science Foundation of China under grant nos. 92270107, 12404568, T2225018, and T2121001, and the Guangzhou Municipal Science and Technology Project (No. 2024A03J0607)
@@ Appendix A Implicit function theorem
where we use $\cdot]_{\bm{\phi}}$ to label the part associated with $\bm{\phi}$} and the corresponding linear coefficients in Eq. 9 and  Eq. 10 should be revised as:
@@ Appendix B Four points movement
Figure 13: The phase difference $\Delta\phi_{\mathrm{implicit}}$ obtained from the solution of the linear equation and $\Delta\phi_{\mathrm{expand}}$ from the series expansion in Eq. 11. The label n-n denotes the number of series used.
@@ Appendix D Geometric Transformation
Appendix E Atoms decay along movement
To investigate heating effects arising from intensity fluctuations during optical trap movement, we developed a one-dimensional semi-classical simulation. The model quantifies the impact of intensity cross-talk that occurs when projecting a trap with a device such as a SLM. The trap’s movement is modeled as a discrete step where the potential from the initial trap ($E_{\rm last}$) exponentially decays while the potential of the final trap ($E_{\rm new}$) increases. The total time-dependent potential, $U(t)$, experienced by a single atom is described by the coherent sum of the two electric fields:
$U(t)=|E_{\rm last}{\rm exp}(-t/\tau)+E_{\rm new}(1-{\rm exp}(-t/\tau))|^{2}$ (15)
The parameter $\tau$ represents the characteristic transition time and was set to 1 ms to reflect the $1~\rm kHz$ frame rate of a state-of-the-art SLM.
Figure 18: Left: Initial and final energy distributions after displacement. Center: Initial vs. final energy of each atom. Right: Position difference distribution after displacement. Atoms with a final energy $\geq 70~\rm\mu K$ or a position difference $\geq 22~\rm\mu m$ are classified as escaped atoms
The simulation was parameterized for a single Rb atom with an initial temperature of $10~\rm\mu K$, loaded into a trap with a $1~\rm\mu m$ waist and a maximum depth of $0.1~\rm mK$. A Monte Carlo method was employed, where the initial position and velocity of the atom were sampled from a Maxwell-Boltzmann distribution for each of 300 simulation runs. The results, summarized in  Fig. 18, demonstrate a slight increase in the average final energy of the atomic ensemble. Crucially, the calculated atom survival probability remains high, indicating that heating from this translation process is minimal and does not significantly compromise atom confinement.